OpenStudy (anonymous):

Simplify the following expression. 2x - 2y + 5z - 2x - y + 3z A.) x - 3y + 8z b.) 2x - 2y + 8z c.) 3y + 8z d.) -3y + 8z

OpenStudy (anonymous):

im not good with this whole problem. i tried solving it again,and again and im fruterared because i have no idea.

OpenStudy (anonymous):

the first thing you should try is to put factors with common variables next to each other so 2x - 2y + 5z - 2x - y + 3z would become 2x -2x -2y -y +5z +3z this makes the problem easier to look at

OpenStudy (anonymous):

next, combine factors with common variables. 2x -2x -2y -y +5z +3z the 2x -2x = 0 the -2y -y = -3y and the 5z +3z = 8z so your answer is 0 -3y +8z or just -3y+8z